### v4

# A PBN-RL-XAI Framework for Discovering a “Hit-and-Run” Therapeutic Strategy in Melanoma

Abstract
Innate resistance to anti-PD-1 immunotherapy remains a major clinical challenge in metastatic melanoma, with the underlying molecular networks being poorly understood. To address this, we constructed a dynamic Probabilistic Boolean Network model using transcriptomic data from patient tumor biopsies to elucidate the regulatory logic governing therapy response. We then employed a reinforcement learning agent to systematically discover optimal, multi-step therapeutic interventions and used explainable artificial intelligence to mechanistically interpret the agent’s control policy. The analysis revealed that a precisely timed, 4-step temporary inhibition of the lysyl oxidase like 2 protein (LOXL2) was the most effective strategy. Our explainable analysis showed that this “hit-and-run” intervention is sufficient to erase the molecular signature driving resistance, allowing the network to self-correct without requiring sustained intervention. This study presents a novel, time-dependent therapeutic hypothesis for overcoming immunotherapy resistance and provides a powerful computational framework for identifying non-obvious intervention protocols in complex biological systems.
Index Terms:
Probabilistic Boolean Networks, Melanoma, Immunotherapy Resistance, Systems Biology, Explainable AI.
I Introduction
Immune checkpoint blockade targeting the programmed death-1 (PD-1) pathway has revolutionized treatment for metastatic melanoma, yet innate resistance affects 60-70% of patients [1]. Unlike acquired resistance that develops during treatment, innate resistance manifests as immediate therapeutic failure, suggesting pre-existing molecular networks that render tumors refractory to immune activation from the outset. The mechanistic basis for this variation in response patterns remains poorly explained, representing a critical unmet clinical need.
Recent transcriptomic analyses have identified gene expression signatures associated with resistance, most notably the Innate anti-PD-1 Resistance (IPRES) signature, which encompasses genes involved in mesenchymal transition, extracellular matrix remodeling, angiogenesis, and wound healing [1]. However, the regulatory logic governing how these genes interact dynamically to establish and maintain resistant phenotypes remains largely opaque. Traditional reductionist approaches focusing on individual biomarkers have failed to capture the systems-level properties that govern therapeutic response, largely due to the inherent complexity of the underlying gene regulatory networks (GRNs) [2].
The challenge of understanding resistance networks is compounded by their dynamic and stochastic nature. Gene regulatory networks exhibit complex emergent behaviors, including multistability, where identical genetic backgrounds can give rise to distinct cellular phenotypes depending on the network’s dynamical state. These systems can become trapped in stable attractor states corresponding to different cellular phenotypes, such as therapy resistance [3]. Identifying effective therapeutic interventions requires understanding not only the static architecture of these networks but also their dynamic properties and control mechanisms.
Current therapeutic discovery approaches face fundamental limitations when applied to complex resistance networks. The vast combinatorial space of possible multi-gene interventions makes exhaustive experimental screening prohibitive. Moreover, optimal intervention strategies may involve precisely timed, transient perturbations rather than sustained pathway inhibition [4], and the stochastic nature of biological networks necessitates probabilistic control strategies.
Probabilistic Boolean Networks (PBNs) offer a powerful framework for modeling gene regulatory systems, combining computational tractability with biological realism through stochastic dynamics [5]. PBNs have successfully modeled various biological phenomena, including cell cycle regulation and cancer progression [6]. Concurrently, advances in reinforcement learning (RL) have opened new avenues for discovering optimal control strategies in complex dynamic systems [7]. Unlike traditional optimization approaches requiring explicit system knowledge, RL agents learn optimal policies through environment interaction, showing promise in biological control applications [8].
Here, we present a novel computational framework integrating PBN modeling with reinforcement learning to systematically discover optimal therapeutic interventions for overcoming innate anti-PD-1 resistance in melanoma. We constructed dynamic PBN models from patient transcriptomic data to capture the regulatory logic governing therapy response, then employed RL agents to identify multi-step intervention strategies. Using explainable artificial intelligence, we mechanistically interpreted the learned control policies to understand their biological basis.
Our analysis reveals that innate resistance corresponds to a fundamental system-level transition from a plastic, multi-stable network state to a rigid, canalized landscape dominated by a JUN/LOXL2 regulatory axis. Most significantly, our RL approach discovered a non-obvious ”hit-and-run” therapeutic strategy: a precisely timed, 4-step temporary inhibition of LOXL2 achieving 93.45% success rate in silico. Explainable AI analysis demonstrated that this transient intervention effectively erases the molecular vulnerability signature driving resistance, allowing the network’s intrinsic dynamics to complete the therapeutic transition without sustained intervention.
This work provides the first application of integrated PBN-RL-explainable AI(XAI) methodology to cancer immunotherapy resistance, offering a generalizable framework for discovering time-dependent therapeutic strategies in complex biological systems. Our work is organized as follows. Section II details our methodology, from data acquisition and network inference to the reinforcement learning and XAI frameworks. Section III presents the results, including the characterization of the resistant network and the discovery and validation of optimal control strategies. Finally, Section IV discusses the implications of our findings and concludes the paper.
II Methods
II-A Data Acquisition and Binarization
We utilized the landmark RNA-sequencing dataset from Hugo et al. [1], publicly available from the Gene Expression Omnibus (accession: GSE78220). Pre-treatment tumor biopsies were stratified into “Responders” (n=15), defined as patients with a Complete or Partial Response, and “Non-Responders” (n=12), defined as patients with Progressive Disease. One on-treatment sample was excluded to ensure the analysis focused exclusively on the innate resistance state. For biological interpretation, gene identifiers in the raw data were mapped to official HUGO Gene Symbols.
To prepare the data for Boolean modeling, the continuous gene expression values were converted into discrete binary states of 0 (OFF/low activity) or 1 (ON’/high activity). The analysis began with the raw count matrix, which was first filtered to include only the core genes. To stabilize the variance across the range of mean expression values, a Variance Stabilizing Transformation (VST)[9] was applied to the filtered count data using the DESeq2 R package[10]. The resulting normalized matrix was then separated into two sub-matrices, one for each cohort.
For each gene within each cohort’s VST-normalized matrix, we employed a Gaussian Mixture Model (GMM)[11] to determine a data-driven binarization threshold. This model-based approach was chosen over simpler methods like a median split because it provides a more principled cutoff. GMM assumes that the observed expression values are generated from two underlying distributions representing ’low’ and ’high’ activity states, and it mathematically identifies the threshold that best separates these putative states rather than imposing an arbitrary split.
Mathematically, the probability density of an expression value $x$ for a given gene is modeled as a weighted sum of two Gaussian components:
$p(x)=\pi_{1}\mathcal{N}(x|\mu_{\text{low}},\sigma^{2})+\pi_{2}\mathcal{N}(x|\mu_{\text{high}},\sigma^{2})$ (1)
where $\pi_{1}$ and $\pi_{2}$ are the mixing coefficients that sum to 1, $\mathcal{N}(x|\mu,\sigma^{2})$ is the Gaussian probability density function, and $(\mu_{\text{low}},\mu_{\text{high}},\sigma^{2})$ are the means and common variance of the two components, respectively. These parameters were fitted to the expression data for each gene using the mclust R package. The binarization threshold $T$ was then defined as the midpoint between the means of the two fitted components, $T=(\mu_{\text{low}}+\mu_{\text{high}})/2$. Samples with an expression value greater than this threshold were assigned a state of 1, while those with values less than or equal to the threshold were assigned a state of 0. This process yielded two final binarized matrices that served as the direct inputs for the subsequent network inference.
II-B Introduction of PBN
To capture the complex and stochastic nature of gene regulatory networks (GRNs), we employ the Probabilistic Boolean Network (PBN) framework [5]. A PBN is a discrete-time dynamical system that models a set of genes as binary variables, where the state of each gene (ON=1, OFF=0) is updated at each time step according to a set of probabilistic rules.
A PBN is formally defined by a set of $n$ genes, $G=\{g_{1},g_{2},\ldots,g_{n}\}$. The state of the network at time $t$ is given by a binary vector $\mathbf{x}(t)=(x_{1}(t),x_{2}(t),\ldots,x_{n}(t))$, where $x_{i}(t)\in\{0,1\}$ is the state of gene $g_{i}$.
For each gene $g_{i}$, there exists a set of $k_{i}$ possible Boolean predictor functions, $F_{i}=\{f_{1}^{(i)},f_{2}^{(i)},\ldots,f_{k_{i}}^{(i)}\}$. Each function $f_{j}^{(i)}:\{0,1\}^{n}\to\{0,1\}$ determines the next state of gene $g_{i}$ based on the current state of all genes in the network. At each time step $t$, a single function $f_{j}^{(i)}$ is selected from the set $F_{i}$ with a corresponding probability $c_{j}^{(i)}$, where:
$P(f=f_{j}^{(i)})=c_{j}^{(i)}\quad\text{and}\quad\sum_{j=1}^{k_{i}}c_{j}^{(i)}=1$ (2)
The state of gene $g_{i}$ at the next time step, $t+1$, is then determined by the selected function: $x_{i}(t+1)=f_{j}^{(i)}(\mathbf{x}(t))$. The selection of a predictor function for each gene is an independent event, defining the transition probabilities between the $2^{n}$ possible states of the network.
The long-term behavior of a PBN is characterized by its attractors. An attractor is a set of states from which the network cannot escape once entered. Mathematically, a non-empty set of states $A$ is an attractor if it is a minimal set for which the probability of transitioning from any state $\mathbf{x}\in A$ to any state $\mathbf{y}\notin A$ is zero. These attractors can be a single steady state (a point attractor) or a set of states visited cyclically (a cycle attractor). In the context of systems biology, attractors represent the stable, long-term phenotypes of the cell, such as proliferation, apoptosis, or in our case, a therapy-resistant state[12].
To quantify the regulatory impact of one gene on another within the network, we utilize the influence of gene $g_{j}$ on gene $g_{i}$. This influence is based on the concept of the sensitivity of the Boolean predictor functions. The influence of $g_{j}$ on $g_{i}$ with respect to a specific function $f_{k}^{(i)}$ is the probability that flipping the state of $g_{j}$ will flip the output of $f_{k}^{(i)}$. The total influence, $I(g_{j}\to g_{i})$, is the expected sensitivity, averaged over all possible predictor functions for gene $g_{i}$:
$I(g_{j}\to g_{i})=\sum_{k=1}^{k_{i}}c_{k}^{(i)}\cdot S(f_{k}^{(i)},g_{j})$ (3)
where $S(f_{k}^{(i)},g_{j})$ is the sensitivity of function $f_{k}^{(i)}$ to gene $g_{j}$. The influence matrix, composed of these values, provides a quantitative map of the network’s regulatory structure.
II-C PKN-constrained PBN construction
The inference of Probabilistic Boolean Networks (PBNs) from data is an NP-hard problem, where the search space for regulatory logic grows double-exponentially with the number of inputs to a given gene, making the analysis of large-scale networks computationally infeasible. Therefore, a critical preliminary step is to constrain the problem space by identifying a core set of the most biologically relevant and dynamically active genes. To this end, we implemented a multi-stage procedure to construct a high-confidence, context-specific Prior Knowledge Network (PKN) that serves as the basis for our tractable PBN model.
II-C1 Initial network foundation
The process began with a broad, literature-based network that was systematically refined into a compact, data-driven core network. The initial network foundation was based on the Kyoto Encyclopedia of Genes and Genomes (KEGG) pathway hsa05235, ”PD-L1 expression and PD-1 checkpoint pathway in cancer,” chosen for its direct relevance to the therapeutic mechanism under investigation. The pathway data was obtained in KGML format, and all constituent gene-gene interactions were parsed and mapped to their corresponding HUGO Gene Symbols.
II-C2 Hybrid gene selection strategy
To distill this foundational network into a focused set of key regulators, we devised a hybrid selection strategy combining mandatory inclusion of expert-defined markers with a data-driven ranking of all other genes. This approach guaranteed the inclusion of five core genes from the Innate anti-PD-1 Resistance (IPRES) signature: AXL, ROR2, WNT5A, LOXL2, and TAGLN. The remaining positions in the core gene list were filled by ranking genes based on a composite score derived from two metrics.
The first metric, a Differential Expression Score ($S_{DE}$), was calculated as the negative base-10 logarithm of the Benjamini-Hochberg adjusted p-value ($-\log_{10}(p_{adj})$)[13] from a differential expression analysis[14] between the responder and non-responder cohorts. The second metric, a Network Influence Score ($S_{Net}$), was calculated from a data-driven gene regulatory network inferred by dynGENIE3[15], capturing a gene’s overall importance as both a regulator and a target.
For the final PKN assembly, both the $S_{DE}$ and $S_{Net}$ scores were min-max normalized to a [0, 1] scale. The final data-driven score for each gene was calculated using the equally weighted formula:
$S_{final}=(0.5\cdot S_{DE,norm})+(0.5\cdot S_{Net,norm})$ (4)
The final core gene list was created by taking the union of the five mandatory IPRES genes and the top 12 genes from this data-driven ranking. The final PKN was then constructed by extracting the subnetwork from the dynGENIE3-inferred weight matrix corresponding only to the interactions among these selected core genes.
II-C3 PBN inference methodology
With the binarized data for the Responder and Non-Responder cohorts, we inferred two distinct PBN models. The inference strategy was designed to identify the most informative regulatory logic for each gene while mitigating the risk of overfitting the model to noise in the data.
The process began by constraining the search space for potential regulators for each target gene. This is a critical step for computational tractability, as testing all possible gene combinations is infeasible. We used the global gene-gene interaction rankings previously generated by dynGENIE3 as a data-driven prior. For each target gene, we selected its top-ranked regulators from this list to form a candidate pool, limiting the pool size to a maximum of 10 genes to balance biological comprehensiveness with computational efficiency.
A key challenge in network inference is selecting the correct number of input regulators, $k$, for each gene’s predictive function. A model that is too simple (small $k$) may underfit and miss critical interactions, while one that is too complex (large $k$) may overfit by modeling noise. To navigate this trade-off, we implemented a data-driven selection criterion based on maximizing the Mutual Information (MI). For each target gene, our algorithm iteratively tested regulator sets of increasing size (up to a maximum of $k_{max}=4$). At each value of $k$, the algorithm identified the single best-predicting Boolean function from all $2^{2^{k}}$ possibilities and calculated its MI score with the target gene’s observed data. The optimal number of regulators, $k^{*}$, was chosen as the value of $k$ that yielded the function with the maximum MI score.
With an optimal regulator set of size $k^{*}$ identified for a given target gene, its probabilistic Boolean update rule was defined. To account for biological stochasticity and measurement noise, we constructed a probabilistic ensemble of the most likely functions rather than selecting only the single best predictor. All $2^{2^{k^{*}}}$ possible Boolean functions were scored using their normalized Mutual Information (MI)[16] value, and the top $N=4$ functions were selected to form the function set $\{f_{1},f_{2},\ldots,f_{N}\}$ for that gene. The probability $P(f_{i})$ of each function was then assigned proportionally to its MI score, as follows:
$P(f_{i})=\frac{MI(f_{i})}{\sum_{j=1}^{N}MI(f_{j})}$ (5)
where $MI(f_{i})$ is the Mutual Information score of function $f_{i}$. An alternative ranking approach using the Coefficient of Determination (CoD)[17] was considered; however, it was found to be impractical as it produced negative values for some functions in our dataset, making MI a more robust metric for this application.
This entire procedure was performed independently for every gene in both the responder and non-responder datasets, yielding the two final, context-specific PBN models used for all subsequent analysis.
II-D Reinforcement Learning for Optimal Dynamic Control
To discover an optimal intervention strategy in a complex biological network, we framed the control of the resistant-state PBN as a reinforcement learning (RL) problem. The state space of the PBN is vast ($2^{n}$ states), rendering methods based on exhaustive search computationally intractable. Furthermore, the inherent stochasticity of the network means a fixed sequence of interventions does not guarantee a specific outcome. RL is well-suited for finding optimal policies in such large, stochastic environments. To formulate the control problem, we utilized the gym-PBN software package, a toolkit designed for the analysis and RL control of Probabilistic Boolean Networks within the Gymnasium framework [18].
We formally model the control problem as a Markov Decision Process (MDP) [19]. The state space, $\mathcal{S}$, comprises the $2^{n}$ possible binary states, $\mathbf{x}$, of the PBN. The action space, $\mathcal{A}$, is a set of discrete actions including “Do Nothing” and flipping the state of any single gene. The transition probabilities, $P(\mathbf{x}_{t+1}|\mathbf{x}_{t},a_{t})$, are implicitly defined by the PBN’s update rules under a given action. To model the clinical time constraint, we set the maximum steps to be 15. Finally, the reward function, $R$, was designed to guide the agent: a large positive reward ($+100$) is given for reaching the target sensitive state, and a penalty ($-5$) for ending an episode in a known resistant attractor. The action cost is set to 00 so that the agent is not prevented from imposing actions during the 15 steps horizon. The agent’s goal is to learn a policy, $\pi(a_{t}|\mathbf{x}_{t})$, that maximizes the expected cumulative reward.
We chose Proximal Policy Optimization (PPO) as our learning algorithm due to its state-of-the-art performance and stability [20]. PPO optimizes the clipped surrogate objective:
$L^{CLIP}(\theta)=\mathbb{E}_{t}\left[\min\left(r_{t}(\theta)A_{t},\text{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)A_{t}\right)\right]$ (6)
where $r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}|\mathbf{x}_{t})}{\pi_{\theta_{old}}(a_{t}|\mathbf{x}_{t})}$ is the probability ratio between new and old policies, $A_{t}$ is the advantage estimate, and $\epsilon$ is the clipping parameter. This clipping mechanism prevents large, destabilizing policy updates by constraining $r_{t}(\theta)\in[1-\epsilon,1+\epsilon]$. We utilized the PPO agent from Stable-baselines3 [21] with default hyperparameters ($\gamma=0.99$, $\lambda=0.95$, $\epsilon=0.2$) and trained for 500,000 timesteps.
II-E Sequential Intervention and Explainable AI Analysis
II-E1 Optimal “Hit-and-Run” Strategy Identification
To identify a more clinically translatable strategy, we also sought to find an optimal transient intervention. We designed a sequential experiment to test a “hit-and-run” therapeutic model, where the rationale is that a brief, initial perturbation may be sufficient to destabilize the resistant state, allowing the network’s own dynamics to resolve towards a sensitive phenotype without sustained intervention. This experiment was conducted in two phases. First, in the priming phase, we simulated a temporary inhibition of a key target gene (LOXL2, MAPK3, or JUN) by clamping its state to 0 for a specific duration, ranging from 1 to 5 simulation steps. Second, in the control phase, the pre-trained PPO agent was given control of the network from its now-primed state. The agent’s success in steering the network to the target sensitive state within the remainder of the 15-step episode was recorded to identify the most effective transient intervention.
II-E2 Mechanistic Policy Explanation via SHAP
To gain biological insight into our agent’s decision-making strategy, we employed SHAP (SHapley Additive exPlanations) [22], a game-theoretic explainable AI method that quantifies each feature’s contribution to model predictions by computing average marginal contributions across all possible feature combinations. We applied SHAP to analyze the agent’s policy in two key scenarios: (1) identifying the context-dependent ”vulnerability signature” of genes that triggered specific interventions, and (2) assessing how transient priming interventions altered this signature. Additionally, we generated SHAP trajectory visualizations to track the evolution of feature importance over entire control episodes, enabling dynamic interpretation of the agent’s reasoning process.
III Results and Discussions
III-A The Resistant Network is Characterized by a Rigid and Canalized Attractor Landscape
To investigate the dynamic consequences of network rewiring, we performed a computational attractor analysis on both the responder and non-responder PBN models using the optPBN MATLAB toolbox[23]. For each PBN, we estimated the attractor landscape by generating an ensemble of 1,000,000 individual Boolean networks (BNs), with each BN stochastically sampled to represent a specific realization of the probabilistic model. We then identified all steady-state and cyclical attractors for each sampled BN. By aggregating these results across all simulations, we calculated the probability of observing each unique attractor, yielding a quantitative map of the system’s dynamics.
The analysis reveals a reorganization of the network’s dynamic landscape, shifting from a plastic, multi-stable state in responders to a rigid, canalized state in non-responders. The responder network exhibits high plasticity, characterized by 22 distinct attractors, with the most probable one accounting for only 12.55% of the total probability basin. In stark contrast, the non-responder network is highly rigid, collapsing into only 16 attractors. A single singleton attractor dominates this landscape, accounting for 50.02% of the total probability, which indicates the network has been rewired to robustly maintain a specific resistant phenotype.
TABLE I: Key cellular phenotypes identified through attractor analysis. Each state is defined by the binary expression (OFF=0, ON=1) of the core IPRES genes AXL, WNT5A, and ROR2, distinguishing between therapy-sensitive and resistant subtypes.
Phenotype Key Gene States
(AXL, WNT5A, ROR2)
Responder (Sensitive) (0, 0, 1)
Resistant (WNT-Driven) (0, 1, 1)
Resistant (AXL-Driven) (1, 0, 1)
Dominant Resistant State (0, 0, 0)
Decoding the dominant attractors provides insight into the specific gene expression patterns that define the sensitive and resistant states. While the responder landscape was fragmented across multiple phenotypes, the analysis of the non-responder landscape revealed that the dominant attractor did not fully match the expected Innate anti-PD-1 Resistance (IPRES) signature (e.g., AXL=1, WNT5A=1, ROR2=1). The key identified cellular phenotypes are summarized in Table I.
This analysis demonstrates that innate resistance corresponds to a fundamental system-level transition to a new, highly stable dynamic regime, where the network becomes hard-wired to occupy a dominant, therapy-resistant attractor.
III-B Network Rewiring Establishes a JUN/LOXL2 Regulatory Axis
Having established that the resistant state is defined by a rigid and dominant attractor, we next sought to identify the specific regulatory changes responsible for this landscape collapse. We performed a comparative analysis of the PBN models, combining a quantitative assessment of gene influence with a detailed inspection of the inferred Boolean logic.
To identify the key regulatory hubs, we calculated a total influence score for each gene by summing all its incoming and outgoing edge weights from the network’s influence matrix. This score quantifies a gene’s overall impact as both a regulator and a target within the system. The analysis revealed a significant shift in the network’s command structure, establishing the transcription factor JUN as the unequivocal master regulator of the resistant state, with LOXL2 emerging as its primary executive arm. As shown in Figure 1, JUN’s total influence increases dramatically in the non-responder network, and LOXL2 exhibits the second-largest gain. This ascent pinpoints the JUN/LOXL2 axis as the central driver of the switch to the therapy-resistant phenotype.
Figure 1: Shift in total gene influence from the responder (blue) to the non-responder (red) model. The plot highlights the change in influence score for each gene, sorted by their final influence in the non-responder state. The significant increase in JUN’s influence, along with notable shifts in LOXL2, illustrates the systemic rewiring of regulatory control.
This dramatic rise in influence is the direct result of a fundamental rewiring of both genes’ regulatory logic. In the sensitive responder network, JUN’s activation is highly conditional, relying on a specific context of present and absent upstream signals. In stark contrast, the logic in the resistant network becomes decisively canalized, collapsing to a simple, robust AND gate that makes JUN’s activation an “all-or-nothing” event: JUN = MAP2K3 AND NRAS AND RELA AND LOXL2 This rewiring creates what is effectively a hard-wired activation switch for the master regulator. Concurrently, the regulatory logic for LOXL2, the axis’s executive arm, undergoes a similar canalization. In the responder network, LOXL2’s state is determined by a complex, probabilistic mixture of functions. In the resistant network, this complexity collapses into a single, deterministic rule that enslaves LOXL2 to its new master regulator: LOXL2 = JUN AND MAP2K3
This hierarchical simplification, where JUN is rewired into a master switch and LOXL2 is rewired into its direct effector, hard-wires the JUN/LOXL2 axis and locks the system into a robustly resistant phenotype.
III-C Reinforcement Learning Identifies the MAP2K3/LOXL2 Axis as the Optimal Dynamic Control Target
To identify an optimal, unconstrained control policy, we trained a reinforcement learning agent to steer the network from three resistant attractor states to the target sensitive phenotype as in Table I. The resulting policy was exceptionally robust, achieving a 100% success rate over 10,000 evaluation trials.
An analysis of the agent’s actions reveals a highly specific and hierarchical control strategy (Table II). The agent’s policy converged on targeting the MAP2K3/LOXL2 axis; its most frequent intervention was to flip MAP2K3, followed by flipping LOXL2. This demonstrates that while static network analysis pointed to JUN as a master regulator, the dynamic, goal-oriented learning process discovered a more efficient control pathway centered on the upstream kinase MAP2K3.
TABLE II: Action frequencies of the trained reinforcement learning agent over 10,000 evaluation episodes. The data reveals a highly specific control policy that converges on flipping MAP2K3/LOXL2 as the primary intervention, identifying it as the optimal dynamic control target.
Action Times Chosen
Flip Gene ’MAP2K3’ 9037
Flip Gene ’LOXL2’ 3366
Flip Gene ’TAGLN’ 659
Flip Gene ’RELA’ 483
Flip Gene ’WNT5A’ 293
Flip Gene ’JUN’ 183
III-D Temporal Inhibition Reveals a “Hit-and-Run” Therapeutic Strategy
To identify a clinically relevant, transient intervention strategy, we performed a comprehensive analysis simulating a temporary “priming” inhibition of key targets (JUN, LOXL2, and MAPK3) for varying durations. After each priming phase, a pre-trained RL agent was tasked with guiding the network to the therapy-sensitive state.
The results, visualized in Fig. 2, identified a highly effective and specific optimal strategy: a 4-step temporary inhibition of LOXL2, which achieved the highest success rate of 93.45%. The non-monotonic relationship between inhibition duration and success rate for both LOXL2 and JUN provides strong evidence for a “hit-and-run” mechanism, where a precisely timed intervention is more effective than a sustained one. A striking and consistent finding across all experiments was that the agent’s overwhelming post-priming policy was to “Do Nothing”. This validates our hypothesis that a brief, targeted intervention is sufficient to destabilize the resistant network, allowing the system’s own dynamics to complete the transition to a sensitive state.
Figure 2: Efficacy of “hit-and-run” priming strategies. The plot shows the control success rate as a function of the inhibition duration for three key targets. A 4-step inhibition of LOXL2 emerges as the optimal transient strategy, significantly outperforming other interventions.
Figure 3: Comparative SHAP analysis for the agent’s decision to flip MAP2K3 versus JUN from a resistant state. The plot shows the SHAP values for each gene’s contribution to the respective action’s logit, revealing a context-dependent policy.
III-E Mechanistic Policy Explanation via SHAP
To deconstruct the sophisticated policy learned by the reinforcement learning agent, we employed SHAP[22]. This allowed us to move beyond observing the agent’s actions to understanding its underlying reasoning, providing mechanistic explanations for its control strategy.
III-E1 Context-Dependent Intervention Choice
We sought to understand why the agent prioritizes flipping MAP2K3 over other influential genes like JUN. A comparative SHAP analysis reveals a highly context-dependent policy (Fig. 3). While high LOXL2 states push the agent towards flipping either MAP2K3 or JUN, the states of NRAS and TAGLN act as decisive signals: high NRAS encourages MAP2K3 targeting, while high TAGLN favors JUN intervention.
This context-dependent strategy aligns with melanoma signaling biology. NRAS, mutated in 15-20% of melanomas [24], activates the MAPK cascade through MAP2K3 [25]. The agent’s preference for MAP2K3 in NRAS-high contexts reflects understanding of this upstream-downstream relationship, intercepting oncogenic signals at a strategic chokepoint. Conversely, TAGLN drives cytoskeletal remodeling and epithelial-mesenchymal transition [26], prompting the agent to target JUN—a key AP-1 transcription factor [27]—to disrupt transcriptional programs driving resistance.
This demonstrates that the agent discerns specific molecular contexts rather than identifying a generic resistant state, selecting the most effective pathway-specific intervention and mirroring precision oncology strategies [28].
III-E2 Validation of the “Hit-and-Run” Mechanism
Next, we used SHAP to explain why the 4-step inhibition of LOXL2 was more successful than the sub-optimal 2-step inhibition of JUN. We analyzed the network state after each priming strategy by measuring the average SHAP importance of each gene. The results show that the optimal LOXL2 priming is more effective because it does a better job of erasing the “vulnerability signature” that triggers the agent’s desire to intervene (Fig. 4). After the LOXL2 inhibition, the residual SHAP values for key driver genes like NRAS and RELA are significantly lower than after the JUN inhibition, indicating the network has been pushed to a more stable, less-resistant state.
Figure 4: Comparison of residual SHAP feature importance after the optimal (4-step LOXL2) versus a sub-optimal (2-step JUN) priming strategy. The more effective LOXL2 priming leaves a state with lower overall SHAP importance, requiring less subsequent intervention.
[…]
Figure 5: SHAP trajectory analysis explaining the agent’s policy. The most frequent action was consistently “Do Nothing,” represented by a dash (-) on the x-axis. Persistent negative SHAP values for NRAS and RELA reveal a vulnerability signature that the agent learns to ignore, validating the “hit-and-run” strategy.
III-E3 Dynamic Visualization of the Control Strategy
Finally, to visualize the agent’s reasoning, we analyzed the average of 10,000 simulated 15-step control episodes (Fig. 5). The resulting SHAP trajectory heatmap reveals that the most frequent action at every step was “Do Nothing,” providing strong quantitative support for a “hit-and-run” control mechanism where the network self-corrects after an initial priming.
The heatmap illustrates a persistent “vulnerability signature” that drives the agent’s policy. At the initial step, we observe a powerful negative SHAP contribution from the key driver genes NRAS and RELA, indicating a strong initial impetus for intervention. While this signal weakens after the first step, it does not disappear; instead, it persists at a consistent, lower intensity for the remainder of the episode. This suggests that even as the network evolves, these genes continue to exert a background pressure that the agent has learned to correctly ignore. Other genes, including LOXL2 and MAPK3, contribute weaker, transient negative signals in the early steps before becoming neutral.
This temporal pattern of network self-organization following an initial perturbation aligns with emerging therapeutic paradigms in cancer treatment. The concept parallels successful clinical strategies such as induction chemotherapy in acute leukemias, where intensive initial treatment creates system instability that resolves into remission [34], and immune checkpoint blockade, where brief T-cell activation triggers sustained anti-tumor responses [35]. The identification of NRAS and RELA as key, persistent vulnerability drivers reflects established understanding that RAS/MAPK hyperactivation and NF-$\kappa$B inflammatory signaling represent critical resistance mechanisms [36, 1], supporting the biological plausibility of the discovered intervention strategy.
IV Conclusion
This study introduced a novel computational framework integrating PBN modeling, reinforcement learning, and explainable AI to deconstruct and control innate immunotherapy resistance in melanoma. Our analysis revealed that resistance corresponds to a fundamental system-level transition from a plastic, multi-stable network to a rigid, canalized landscape driven by a JUN/LOXL2 regulatory axis that locks the system into a robustly resistant state.
Most significantly, our reinforcement learning approach discovered a non-obvious ”hit-and-run” therapeutic strategy: a precisely timed, 4-step temporary inhibition of LOXL2 achieving over 93% success in silico. Explainable AI analysis demonstrated that this transient intervention erases the molecular vulnerability signature stabilizing resistance, allowing the network’s intrinsic dynamics to complete the therapeutic transition.
Our findings are based on an in silico model with inherent limitations, including its binary gene representation, focus on a core gene set, and the specific reward configuration for reinforcement learning. For instance, using a positive action cost instead of zero would penalize frequent interventions and could yield a different optimal policy. Consequently, the identified strategies should be interpreted as computational hypotheses that require rigorous experimental validation in cell lines and animal models before any clinical consideration.
The discovery of time-dependent optimal strategies challenges traditional paradigms focused on sustained pathway inhibition. The integrated PBN-RL-XAI pipeline offers a generalizable approach for discovering novel control strategies in complex, treatment-resistant diseases, providing a powerful framework for future therapeutic development.
V Code Availability
All source code developed for the PBN inference, reinforcement learning, and subsequent analyses presented in this paper are publicly available in our GitHub repository at https://github.com/Liu-Zhonglin/pbn-melanoma-project.
Acknowledgment
The author would like to express sincere gratitude to Professor Zhang Louxin for his valuable advice on potential improvements to this work. Special thanks are also extended to Professor Ching Wai Ki for his thorough internal review and constructive feedback.

### v5

@@
 This study presents a novel, time-dependent therapeutic hypothesis for overcoming immunotherapy resistance and provides a powerful computational framework for identifying non-obvious intervention protocols in complex biological systems.
-Index Terms:
-Probabilistic Boolean Networks, Melanoma, Immunotherapy Resistance, Systems Biology, Explainable AI.
 I Introduction
@@
 Traditional reductionist approaches focusing on individual biomarkers have failed to capture the systems-level properties that govern therapeutic response, largely due to the inherent complexity of the underlying gene regulatory networks (GRNs) [2].
-The challenge of understanding resistance networks is compounded by their dynamic and stochastic nature.
-Gene regulatory networks exhibit complex emergent behaviors, including multistability, where identical genetic backgrounds can give rise to distinct cellular phenotypes depending on the network’s dynamical state.
-These systems can become trapped in stable attractor states corresponding to different cellular phenotypes, such as therapy resistance [3].
-Identifying effective therapeutic interventions requires understanding not only the static architecture of these networks but also their dynamic properties and control mechanisms.
 Current therapeutic discovery approaches face fundamental limitations when applied to complex resistance networks.
@@
 Moreover, optimal intervention strategies may involve precisely timed, transient perturbations rather than sustained pathway inhibition [3], and the stochastic nature of biological networks necessitates probabilistic control strategies.
-Probabilistic Boolean Networks (PBNs) offer a powerful framework for modeling gene regulatory systems, combining computational tractability with biological realism through stochastic dynamics [5].
-PBNs have successfully modeled various biological phenomena, including cell cycle regulation and cancer progression [6].
-Concurrently, advances in reinforcement learning (RL) have opened new avenues for discovering optimal control strategies in complex dynamic systems [7].
-Unlike traditional optimization approaches requiring explicit system knowledge, RL agents learn optimal policies through environment interaction, showing promise in biological control applications [8].
-Here, we present a novel computational framework integrating PBN modeling with reinforcement learning to systematically discover optimal therapeutic interventions for overcoming innate anti-PD-1 resistance in melanoma.
+While existing computational approaches for therapeutic discovery have explored various methodologies—including Boolean network control [4], dynamical systems optimization [5], and machine learning-based drug repurposing [6]—these methods typically focus on static network analysis or sustained interventions.
+Our approach uniquely integrates Probabilistic Boolean Networks (PBNs) with reinforcement learning (RL) to discover time-dependent, transient intervention strategies.
+PBNs combine computational tractability with biological realism through stochastic dynamics [7], while RL enables discovery of optimal control policies without requiring explicit system knowledge [8].
+This integration allows systematic exploration of dynamic intervention strategies that existing static approaches cannot capture.
+Here, we present this novel computational framework to systematically discover optimal therapeutic interventions for overcoming innate anti-PD-1 resistance in melanoma.
 We constructed dynamic PBN models from patient transcriptomic data to capture the regulatory logic governing therapy response, then employed RL agents to identify multi-step intervention strategies.
-Using explainable artificial intelligence, we mechanistically interpreted the learned control policies to understand their biological basis.
+Using explainable artificial intelligence (XAI), we mechanistically interpreted the learned control policies to understand their biological basis.
 Our analysis reveals that innate resistance corresponds to a fundamental system-level transition from a plastic, multi-stable network state to a rigid, canalized landscape dominated by a JUN/LOXL2 regulatory axis.
-Most significantly, our RL approach discovered a non-obvious ”hit-and-run” therapeutic strategy: a precisely timed, 4-step temporary inhibition of LOXL2 achieving 93.45% success rate in silico.
-Explainable AI analysis demonstrated that this transient intervention effectively erases the molecular vulnerability signature driving resistance, allowing the network’s intrinsic dynamics to complete the therapeutic transition without sustained intervention.
-This work provides the first application of integrated PBN-RL-explainable AI(XAI) methodology to cancer immunotherapy resistance, offering a generalizable framework for discovering time-dependent therapeutic strategies in complex biological systems.
-Our work is organized as follows.
-Section II details our methodology, from data acquisition and network inference to the reinforcement learning and XAI frameworks.
-Section III presents the results, including the characterization of the resistant network and the discovery and validation of optimal control strategies.
-Finally, Section IV discusses the implications of our findings and concludes the paper.
+Most significantly, our RL approach discovered a non-obvious “hit-and-run” therapeutic strategy: a precisely timed, 4-step temporary inhibition of LOXL2 achieving 93.45% success rate in silico.
+This computational hypothesis requires rigorous experimental validation before clinical consideration, yet demonstrates the framework’s potential for discovering non-intuitive therapeutic strategies in complex biological systems.
 II Methods
@@
 We utilized the landmark RNA-sequencing dataset from Hugo et al.
-[1], publicly available from the Gene Expression Omnibus (accession: GSE78220).
-Pre-treatment tumor biopsies were stratified into “Responders” (n=15), defined as patients with a Complete or Partial Response, and “Non-Responders” (n=12), defined as patients with Progressive Disease.
-One on-treatment sample was excluded to ensure the analysis focused exclusively on the innate resistance state.
-For biological interpretation, gene identifiers in the raw data were mapped to official HUGO Gene Symbols.
-To prepare the data for Boolean modeling, the continuous gene expression values were converted into discrete binary states of 0 (OFF/low activity) or 1 (ON’/high activity).
-The analysis began with the raw count matrix, which was first filtered to include only the core genes.
-To stabilize the variance across the range of mean expression values, a Variance Stabilizing Transformation (VST)[9] was applied to the filtered count data using the DESeq2 R package[10].
-The resulting normalized matrix was then separated into two sub-matrices, one for each cohort.
-For each gene within each cohort’s VST-normalized matrix, we employed a Gaussian Mixture Model (GMM)[11] to determine a data-driven binarization threshold.
-This model-based approach was chosen over simpler methods like a median split because it provides a more principled cutoff.
-GMM assumes that the observed expression values are generated from two underlying distributions representing ’low’ and ’high’ activity states, and it mathematically identifies the threshold that best separates these putative states rather than imposing an arbitrary split.
-Mathematically, the probability density of an expression value $x$ for a given gene is modeled as a weighted sum of two Gaussian components:
+[1], publicly available from the Gene Expression Omnibus[9] (accession: GSE78220).
+This dataset comprises transcriptomic profiles from 28 pre-treatment melanoma biopsies obtained from patients undergoing anti-PD-1 checkpoint inhibitor therapy (pembrolizumab or nivolumab) for metastatic melanoma.
+The cohort included patients aged 19-84 years (median 60 years) with 75% male representation and predominantly advanced disease (89% M1b/M1c stage).
+Approximately 50% had prior MAPK inhibitor treatment.
+Pre-treatment tumor biopsies were stratified into “Responders” (n=15), defined as patients achieving Complete or Partial Response, and “Non-Responders” (n=13), defined as patients with Progressive Disease according to immune-related response criteria.
+One on-treatment sample was excluded to ensure analysis focused exclusively on the innate resistance state.
+Gene identifiers were mapped to official HUGO Gene Symbols for biological interpretation.
+To prepare data for Boolean modeling, continuous gene expression values were converted into discrete binary states of 0 (OFF/low activity) or 1 (ON/high activity).
+Raw count data were filtered to include only core genes, then Variance Stabilizing Transformation (VST) [10] was applied using DESeq2 [11] to stabilize variance across expression ranges.
+For each gene within each cohort’s VST-normalized matrix, we employed a Gaussian Mixture Model (GMM) [12] to determine data-driven binarization thresholds.
+The probability density of expression value $x$ was modeled as:
 $p(x)=\pi_{1}\mathcal{N}(x|\mu_{\text{low}},\sigma^{2})+\pi_{2}\mathcal{N}(x|\mu_{\text{high}},\sigma^{2})$ (1)
-where $\pi_{1}$ and $\pi_{2}$ are the mixing coefficients that sum to 1, $\mathcal{N}(x|\mu,\sigma^{2})$ is the Gaussian probability density function, and $(\mu_{\text{low}},\mu_{\text{high}},\sigma^{2})$ are the means and common variance of the two components, respectively.
-These parameters were fitted to the expression data for each gene using the mclust R package.
-The binarization threshold $T$ was then defined as the midpoint between the means of the two fitted components, $T=(\mu_{\text{low}}+\mu_{\text{high}})/2$.
-Samples with an expression value greater than this threshold were assigned a state of 1, while those with values less than or equal to the threshold were assigned a state of 0.
-This process yielded two final binarized matrices that served as the direct inputs for the subsequent network inference.
+where $\pi_{1}$ and $\pi_{2}$ are mixing coefficients summing to 1, and $(\mu_{\text{low}},\mu_{\text{high}},\sigma^{2})$ represent component means and variance.
+Parameters were fitted using mclust R package, with binarization threshold $T=(\mu_{\text{low}}+\mu_{\text{high}})/2$.
+This yielded two binarized matrices for subsequent network inference.
 II-B Introduction of PBN
 To capture the complex and stochastic nature of gene regulatory networks (GRNs), we employ the Probabilistic Boolean Network (PBN) framework [7].
-A PBN is a discrete-time dynamical system that models a set of genes as binary variables, where the state of each gene (ON=1, OFF=0) is updated at each time step according to a set of probabilistic rules.
-A PBN is formally defined by a set of $n$ genes, $G=\{g_{1},g_{2},\ldots,g_{n}\}$.
-The state of the network at time $t$ is given by a binary vector $\mathbf{x}(t)=(x_{1}(t),x_{2}(t),\ldots,x_{n}(t))$, where $x_{i}(t)\in\{0,1\}$ is the state of gene $g_{i}$.
-For each gene $g_{i}$, there exists a set of $k_{i}$ possible Boolean predictor functions, $F_{i}=\{f_{1}^{(i)},f_{2}^{(i)},\ldots,f_{k_{i}}^{(i)}\}$.
-Each function $f_{j}^{(i)}:\{0,1\}^{n}\to\{0,1\}$ determines the next state of gene $g_{i}$ based on the current state of all genes in the network.
-At each time step $t$, a single function $f_{j}^{(i)}$ is selected from the set $F_{i}$ with a corresponding probability $c_{j}^{(i)}$, where:
-$P(f=f_{j}^{(i)})=c_{j}^{(i)}\quad\text{and}\quad\sum_{j=1}^{k_{i}}c_{j}^{(i)}=1$ (2)
-The state of gene $g_{i}$ at the next time step, $t+1$, is then determined by the selected function: $x_{i}(t+1)=f_{j}^{(i)}(\mathbf{x}(t))$.
-The selection of a predictor function for each gene is an independent event, defining the transition probabilities between the $2^{n}$ possible states of the network.
-The long-term behavior of a PBN is characterized by its attractors.
-An attractor is a set of states from which the network cannot escape once entered.
-Mathematically, a non-empty set of states $A$ is an attractor if it is a minimal set for which the probability of transitioning from any state $\mathbf{x}\in A$ to any state $\mathbf{y}\notin A$ is zero.
-These attractors can be a single steady state (a point attractor) or a set of states visited cyclically (a cycle attractor).
-In the context of systems biology, attractors represent the stable, long-term phenotypes of the cell, such as proliferation, apoptosis, or in our case, a therapy-resistant state[12].
-To quantify the regulatory impact of one gene on another within the network, we utilize the influence of gene $g_{j}$ on gene $g_{i}$.
-This influence is based on the concept of the sensitivity of the Boolean predictor functions.
-The influence of $g_{j}$ on $g_{i}$ with respect to a specific function $f_{k}^{(i)}$ is the probability that flipping the state of $g_{j}$ will flip the output of $f_{k}^{(i)}$.
-The total influence, $I(g_{j}\to g_{i})$, is the expected sensitivity, averaged over all possible predictor functions for gene $g_{i}$:
+A PBN models $n$ genes as binary variables with state vector $\mathbf{x}(t)=(x_{1}(t),\ldots,x_{n}(t))$ where $x_{i}(t)\in\{0,1\}$ represents gene $g_{i}$’s state at time $t$.
+For each gene $g_{i}$, there exists a set of $k_{i}$ Boolean predictor functions $F_{i}=\{f_{1}^{(i)},\ldots,f_{k_{i}}^{(i)}\}$.
+At each time step, function $f_{j}^{(i)}$ is selected with probability $c_{j}^{(i)}$ where $\sum_{j=1}^{k_{i}}c_{j}^{(i)}=1$, determining the next state: $x_{i}(t+1)=f_{j}^{(i)}(\mathbf{x}(t))$.
+The long-term behavior is characterized by attractors—minimal sets of states $A$ where $P(\mathbf{x}(t+1)\in A|\mathbf{x}(t)\in A)=1$ and $P(\mathbf{x}(t+1)\notin A|\mathbf{x}(t)\in A)=0$.
+In systems biology, attractors represent stable cellular phenotypes such as therapy-resistant states [13].
+To quantify regulatory relationships, we compute the influence of gene $g_{j}$ on $g_{i}$ as the expected sensitivity across all predictor functions:
 $I(g_{j}\to g_{i})=\sum_{k=1}^{k_{i}}c_{k}^{(i)}\cdot S(f_{k}^{(i)},g_{j})$ (2)
-where $S(f_{k}^{(i)},g_{j})$ is the sensitivity of function $f_{k}^{(i)}$ to gene $g_{j}$.
-The influence matrix, composed of these values, provides a quantitative map of the network’s regulatory structure.
+where $S(f_{k}^{(i)},g_{j})$ is the sensitivity of function $f_{k}^{(i)}$ to gene $g_{j}$. The resulting influence matrix provides a quantitative map of the network’s regulatory structure.
 II-C PKN-constrained PBN construction
-The inference of Probabilistic Boolean Networks (PBNs) from data is an NP-hard problem, where the search space for regulatory logic grows double-exponentially with the number of inputs to a given gene, making the analysis of large-scale networks computationally infeasible.
-Therefore, a critical preliminary step is to constrain the problem space by identifying a core set of the most biologically relevant and dynamically active genes.
-To this end, we implemented a multi-stage procedure to construct a high-confidence, context-specific Prior Knowledge Network (PKN) that serves as the basis for our tractable PBN model.
-II-C1 Initial network foundation
-The process began with a broad, literature-based network that was systematically refined into a compact, data-driven core network.
-The initial network foundation was based on the Kyoto Encyclopedia of Genes and Genomes (KEGG) pathway hsa05235, ”PD-L1 expression and PD-1 checkpoint pathway in cancer,” chosen for its direct relevance to the therapeutic mechanism under investigation.
-The pathway data was obtained in KGML format, and all constituent gene-gene interactions were parsed and mapped to their corresponding HUGO Gene Symbols.
-II-C2 Hybrid gene selection strategy
-To distill this foundational network into a focused set of key regulators, we devised a hybrid selection strategy combining mandatory inclusion of expert-defined markers with a data-driven ranking of all other genes.
-This approach guaranteed the inclusion of five core genes from the Innate anti-PD-1 Resistance (IPRES) signature: AXL, ROR2, WNT5A, LOXL2, and TAGLN.
-The remaining positions in the core gene list were filled by ranking genes based on a composite score derived from two metrics.
-The first metric, a Differential Expression Score ($S_{DE}$), was calculated as the negative base-10 logarithm of the Benjamini-Hochberg adjusted p-value ($-\log_{10}(p_{adj})$)[13] from a differential expression analysis[14] between the responder and non-responder cohorts.
-The second metric, a Network Influence Score ($S_{Net}$), was calculated from a data-driven gene regulatory network inferred by dynGENIE3[15], capturing a gene’s overall importance as both a regulator and a target.
-For the final PKN assembly, both the $S_{DE}$ and $S_{Net}$ scores were min-max normalized to a [0, 1] scale.
-The final data-driven score for each gene was calculated using the equally weighted formula:
-$S_{final}=(0.5\cdot S_{DE,norm})+(0.5\cdot S_{Net,norm})$ (4)
-The final core gene list was created by taking the union of the five mandatory IPRES genes and the top 12 genes from this data-driven ranking.
-The final PKN was then constructed by extracting the subnetwork from the dynGENIE3-inferred weight matrix corresponding only to the interactions among these selected core genes.
-II-C3 PBN inference methodology
-With the binarized data for the Responder and Non-Responder cohorts, we inferred two distinct PBN models.
-The inference strategy was designed to identify the most informative regulatory logic for each gene while mitigating the risk of overfitting the model to noise in the data.
-The process began by constraining the search space for potential regulators for each target gene.
-This is a critical step for computational tractability, as testing all possible gene combinations is infeasible.
-We used the global gene-gene interaction rankings previously generated by dynGENIE3 as a data-driven prior.
-For each target gene, we selected its top-ranked regulators from this list to form a candidate pool, limiting the pool size to a maximum of 10 genes to balance biological comprehensiveness with computational efficiency.
-A key challenge in network inference is selecting the correct number of input regulators, $k$, for each gene’s predictive function.
-A model that is too simple (small $k$) may underfit and miss critical interactions, while one that is too complex (large $k$) may overfit by modeling noise.
-To navigate this trade-off, we implemented a data-driven selection criterion based on maximizing the Mutual Information (MI).
-For each target gene, our algorithm iteratively tested regulator sets of increasing size (up to a maximum of $k_{max}=4$).
-At each value of $k$, the algorithm identified the single best-predicting Boolean function from all $2^{2^{k}}$ possibilities and calculated its MI score with the target gene’s observed data.
-The optimal number of regulators, $k^{*}$, was chosen as the value of $k$ that yielded the function with the maximum MI score.
-With an optimal regulator set of size $k^{*}$ identified for a given target gene, its probabilistic Boolean update rule was defined.
-To account for biological stochasticity and measurement noise, we constructed a probabilistic ensemble of the most likely functions rather than selecting only the single best predictor.
-All $2^{2^{k^{*}}}$ possible Boolean functions were scored using their normalized Mutual Information (MI)[16] value, and the top $N=4$ functions were selected to form the function set $\{f_{1},f_{2},\ldots,f_{N}\}$ for that gene.
-The probability $P(f_{i})$ of each function was then assigned proportionally to its MI score, as follows:
+PBN inference is computationally intractable for large networks due to double-exponential growth in search space.
+We developed a multi-stage procedure to construct a tractable, biologically-relevant 12-gene core network from melanoma anti-PD-1 resistance data.
+We began with the KEGG pathway hsa05235 “PD-L1 expression and PD-1 checkpoint pathway” [14], directly relevant to anti-PD-1 therapy mechanisms.
+Our hybrid selection strategy mandatorily included five IPRES signature genes (AXL, ROR2, WNT5A, LOXL2, TAGLN) [1] known to drive therapy resistance through mesenchymal transition and angiogenesis pathways.
+The remaining seven genes were selected using a composite score combining differential expression significance ($S_{DE}$) calculated as $\log_{10}(p_{adj})$ from Benjamini-Hochberg adjusted p-values [15] from differential expression analysis [16], and network centrality ($S_{Net}$) from dynGENIE3 inference [17]:
+$S_{final}=0.5\cdot S_{DE,norm}+0.5\cdot S_{Net,norm}$ (3)
+The 12-gene limitation balances computational tractability ($2^{12}=4096$ states) with biological comprehensiveness.
+The selected genes span key resistance mechanisms: immune checkpoint regulation (PDCD1, CD274), mesenchymal transition (AXL, ROR2, WNT5A), extracellular matrix remodeling (LOXL2, TAGLN), and metabolic reprogramming.
+Prior studies demonstrate that core regulatory modules of 10-15 genes can capture essential cellular dynamics [7].
+For PBN inference, we constrained potential regulators for each target gene using dynGENIE3 rankings, testing regulator sets of size $k=1$ to $k_{max}=4$.
+The optimal $k^{*}$ was selected by maximizing Mutual Information (MI) [18].
+For each gene’s probabilistic update rule, we selected the top $N=4$ Boolean functions, with probabilities assigned proportionally:
 $P(f_{i})=\frac{MI(f_{i})}{\sum_{j=1}^{N}MI(f_{j})}$ (4)
-where $MI(f_{i})$ is the Mutual Information score of function $f_{i}$.
-An alternative ranking approach using the Coefficient of Determination (CoD)[17] was considered; however, it was found to be impractical as it produced negative values for some functions in our dataset, making MI a more robust metric for this application.
-This entire procedure was performed independently for every gene in both the responder and non-responder datasets, yielding the two final, context-specific PBN models used for all subsequent analysis.
+We also evaluated Coefficient of Determination (CoD) [19] as an alternative metric, but MI proved superior due to CoD’s susceptibility to negative values in our dataset..
+This procedure generated two context-specific PBN models for Responder and Non-Responder cohorts, capturing distinct regulatory logic underlying therapy response patterns.
 II-D Reinforcement Learning for Optimal Dynamic Control
-To discover an optimal intervention strategy in a complex biological network, we framed the control of the resistant-state PBN as a reinforcement learning (RL) problem.
-The state space of the PBN is vast ($2^{n}$ states), rendering methods based on exhaustive search computationally intractable.
-Furthermore, the inherent stochasticity of the network means a fixed sequence of interventions does not guarantee a specific outcome.
-RL is well-suited for finding optimal policies in such large, stochastic environments.
-To formulate the control problem, we utilized the gym-PBN software package, a toolkit designed for the analysis and RL control of Probabilistic Boolean Networks within the Gymnasium framework [18].
-We formally model the control problem as a Markov Decision Process (MDP) [19].
-The state space, $\mathcal{S}$, comprises the $2^{n}$ possible binary states, $\mathbf{x}$, of the PBN.
-The action space, $\mathcal{A}$, is a set of discrete actions including “Do Nothing” and flipping the state of any single gene.
-The transition probabilities, $P(\mathbf{x}_{t+1}|\mathbf{x}_{t},a_{t})$, are implicitly defined by the PBN’s update rules under a given action.
-To model the clinical time constraint, we set the maximum steps to be 15.
-Finally, the reward function, $R$, was designed to guide the agent: a large positive reward ($+100$) is given for reaching the target sensitive state, and a penalty ($-5$) for ending an episode in a known resistant attractor.
-The action cost is set to 00 so that the agent is not prevented from imposing actions during the 15 steps horizon.
-The agent’s goal is to learn a policy, $\pi(a_{t}|\mathbf{x}_{t})$, that maximizes the expected cumulative reward.
-We chose Proximal Policy Optimization (PPO) as our learning algorithm due to its state-of-the-art performance and stability [20]. PPO optimizes the clipped surrogate objective:
+To discover optimal intervention strategies in the resistant-state PBN, we formulated the control problem as a Markov Decision Process (MDP) [20] using the gym-PBN framework [21].
+The state space $\mathcal{S}$ comprises $2^{12}=4096$ binary network states, with action space $\mathcal{A}$ including “Do Nothing” and single-gene state flips.
+Transition probabilities $P(\mathbf{x}_{t+1}|\mathbf{x}_{t},a_{t})$ are defined by PBN update rules under interventions.
+The reward function incentivizes therapeutic transitions: +100+100 for reaching sensitive states, 5-5 for resistant attractors, and 0 action cost to encourage active intervention within the 15-step clinical horizon.
+This design reflects the clinical priority of escaping resistance over minimizing intervention frequency.
+We employed Proximal Policy Optimization (PPO) [22] for its proven stability in complex biological control tasks. PPO optimizes the clipped surrogate objective:
 $L^{CLIP}(\theta)=\mathbb{E}_{t}\left[\min\left(r_{t}(\theta)A_{t},\text{clip}(r_{t}(\theta),1-\epsilon,1+\epsilon)A_{t}\right)\right]$ (5)
-where $r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}|\mathbf{x}_{t})}{\pi_{\theta_{old}}(a_{t}|\mathbf{x}_{t})}$ is the probability ratio between new and old policies, $A_{t}$ is the advantage estimate, and $\epsilon$ is the clipping parameter.
-This clipping mechanism prevents large, destabilizing policy updates by constraining $r_{t}(\theta)\in[1-\epsilon,1+\epsilon]$.
-We utilized the PPO agent from Stable-baselines3 [21] with default hyperparameters ($\gamma=0.99$, $\lambda=0.95$, $\epsilon=0.2$) and trained for 500,000 timesteps.
+where $r_{t}(\theta)=\frac{\pi_{\theta}(a_{t}|\mathbf{x}_{t})}{\pi_{\theta_{old}}(a_{t}|\mathbf{x}_{t})}$ represents the policy ratio and $A_{t}$ the advantage estimate.
+We employed standard PPO hyperparameters from the Stable-Baselines3  [23] implementation: discount factor $\gamma=0.99$ for long-term reward optimization, GAE parameter $\lambda=0.95$ to balance bias-variance in advantage estimation, and clipping parameter $\epsilon=0.2$ to prevent destabilizing policy updates while maintaining exploration.
+These hyperparameters were determined through preliminary experiments to ensure stable performance on our PBN model..
 II-E Sequential Intervention and Explainable AI Analysis
 II-E1 Optimal “Hit-and-Run” Strategy Identification
-To identify a more clinically translatable strategy, we also sought to find an optimal transient intervention.
-We designed a sequential experiment to test a “hit-and-run” therapeutic model, where the rationale is that a brief, initial perturbation may be sufficient to destabilize the resistant state, allowing the network’s own dynamics to resolve towards a sensitive phenotype without sustained intervention.
-This experiment was conducted in two phases.
-First, in the priming phase, we simulated a temporary inhibition of a key target gene (LOXL2, MAPK3, or JUN) by clamping its state to 0 for a specific duration, ranging from 1 to 5 simulation steps.
-Second, in the control phase, the pre-trained PPO agent was given control of the network from its now-primed state.
-The agent’s success in steering the network to the target sensitive state within the remainder of the 15-step episode was recorded to identify the most effective transient intervention.
+We designed a sequential experiment to test clinically translatable transient interventions, where brief perturbations destabilize resistant states without sustained therapy.
+The protocol consisted of two phases: a priming phase with temporary target gene inhibition (LOXL2, MAPK3, or JUN clamped to 0 for 1-5 steps), followed by a control phase where the pre-trained PPO agent guided the primed network toward sensitive states within the remaining 15-step episode.
+Success rates identified optimal transient intervention strategies.
 II-E2 Mechanistic Policy Explanation via SHAP
-To gain biological insight into our agent’s decision-making strategy, we employed SHAP (SHapley Additive exPlanations) [22], a game-theoretic explainable AI method that quantifies each feature’s contribution to model predictions by computing average marginal contributions across all possible feature combinations.
-We applied SHAP to analyze the agent’s policy in two key scenarios: (1) identifying the context-dependent ”vulnerability signature” of genes that triggered specific interventions, and (2) assessing how transient priming interventions altered this signature.
-Additionally, we generated SHAP trajectory visualizations to track the evolution of feature importance over entire control episodes, enabling dynamic interpretation of the agent’s reasoning process.
+We employed SHAP (SHapley Additive exPlanations) [24] to analyze agent decision-making by quantifying each gene’s contribution to intervention choices through game-theoretic marginal contribution analysis.
+SHAP analysis identified context-dependent vulnerability signatures triggering specific interventions and assessed how transient priming altered these signatures.
+Trajectory visualizations tracked feature importance evolution across control episodes for dynamic interpretation of agent reasoning.
 III Results and Discussions
-III-A The Resistant Network is Characterized by a Rigid and Canalized Attractor Landscape
-To investigate the dynamic consequences of network rewiring, we performed a computational attractor analysis on both the responder and non-responder PBN models using the optPBN MATLAB toolbox[23].
-For each PBN, we estimated the attractor landscape by generating an ensemble of 1,000,000 individual Boolean networks (BNs), with each BN stochastically sampled to represent a specific realization of the probabilistic model.
-We then identified all steady-state and cyclical attractors for each sampled BN.
-By aggregating these results across all simulations, we calculated the probability of observing each unique attractor, yielding a quantitative map of the system’s dynamics.
-The analysis reveals a reorganization of the network’s dynamic landscape, shifting from a plastic, multi-stable state in responders to a rigid, canalized state in non-responders.
-The responder network exhibits high plasticity, characterized by 22 distinct attractors, with the most probable one accounting for only 12.55% of the total probability basin.
-In stark contrast, the non-responder network is highly rigid, collapsing into only 16 attractors.
-A single singleton attractor dominates this landscape, accounting for 50.02% of the total probability, which indicates the network has been rewired to robustly maintain a specific resistant phenotype.
-TABLE I: Key cellular phenotypes identified through attractor analysis.
-Each state is defined by the binary expression (OFF=0, ON=1) of the core IPRES genes AXL, WNT5A, and ROR2, distinguishing between therapy-sensitive and resistant subtypes.
-Phenotype Key Gene States
-(AXL, WNT5A, ROR2)
+III-A Network Topology Rewiring Drives Attractor Landscape Rigidification
+To investigate how network rewiring affects system dynamics, we performed computational attractor analysis on both PBN models using the optPBN MATLAB toolbox[25].
+Each PBN was sampled as 1,000,000 individual Boolean networks to map the probability distribution of steady-state and cyclical attractors.
+Network topology comparison reveals fundamental structural differences between responder and non-responder networks (detailed network visualizations in Appendix Fig.
+6).
+The non-responder network exhibits higher connectivity (45 vs 41 edges, density 0.341 vs 0.311) with TAGLN functioning as a resistance hub receiving inputs from RELA and JUN while regulating ROR2.
+Conversely, the responder network shows simplified TAGLN regulation and distinct AXL pathway rewiring toward LOXL2-mediated ECM remodeling.
+Quantitative analysis of node-level topology differences (Appendix Fig.
+7) identifies nodes with substantial changes: absolute in-degree or out-degree differences $\geq$1, or average mutual information differences $\geq$0.2.
+TAGLN emerges as the most dramatically altered node, with 3-fold reduced in-degree connectivity in responders (-3 inputs, -1 output), confirming its critical role as a resistance coordination hub that becomes disrupted in treatment-responsive networks.
+These topological differences drive a dramatic reorganization of the dynamic landscape.
+The responder network exhibits high plasticity with 22 distinct attractors, where the most probable accounts for only 12.55% of the total probability basin.
+In contrast, the non-responder network rigidifies into 16 attractors dominated by a single state capturing 50.02% of probability mass, indicating robust maintenance of resistance through the TAGLN-coordinated regulatory circuit.
+TABLE I: Key Cellular Phenotypes Identified Through Attractor Analysis
+Phenotype (AXL, WNT5A, ROR2)
 Responder (Sensitive) (0, 0, 1)
@@
 Dominant Resistant State (0, 0, 0)
-Decoding the dominant attractors provides insight into the specific gene expression patterns that define the sensitive and resistant states.
-While the responder landscape was fragmented across multiple phenotypes, the analysis of the non-responder landscape revealed that the dominant attractor did not fully match the expected Innate anti-PD-1 Resistance (IPRES) signature (e.g., AXL=1, WNT5A=1, ROR2=1).
-The key identified cellular phenotypes are summarized in Table I.
-This analysis demonstrates that innate resistance corresponds to a fundamental system-level transition to a new, highly stable dynamic regime, where the network becomes hard-wired to occupy a dominant, therapy-resistant attractor.
+This analysis demonstrates that innate resistance represents a system-level transition from plastic, multi-stable dynamics to a rigid, canalized regime where specific regulatory circuits (particularly the TAGLN-ROR2 axis) become hard-wired to maintain resistance.
+The quantitative topology analysis provides mechanistic insight into how network rewiring patterns determine therapeutic response outcomes.
 III-B Network Rewiring Establishes a JUN/LOXL2 Regulatory Axis
@@
 The analysis revealed a significant shift in the network’s command structure, establishing the transcription factor JUN as the unequivocal master regulator of the resistant state, with LOXL2 emerging as its primary executive arm.
-As shown in Figure 1, JUN’s total influence increases dramatically in the non-responder network, and LOXL2 exhibits the second-largest gain.
+As shown in Fig.
+1, JUN’s total influence increases dramatically in the non-responder network, and LOXL2 exhibits the second-largest gain.
 This ascent pinpoints the JUN/LOXL2 axis as the central driver of the switch to the therapy-resistant phenotype.
@@
 The plot highlights the change in influence score for each gene, sorted by their final influence in the non-responder state.
-The significant increase in JUN’s influence, along with notable shifts in LOXL2, illustrates the systemic rewiring of regulatory control.
 This dramatic rise in influence is the direct result of a fundamental rewiring of both genes’ regulatory logic.
@@
 This hierarchical simplification, where JUN is rewired into a master switch and LOXL2 is rewired into its direct effector, hard-wires the JUN/LOXL2 axis and locks the system into a robustly resistant phenotype.
-III-C Reinforcement Learning Identifies the MAP2K3/LOXL2 Axis as the Optimal Dynamic Control Target
-To identify an optimal, unconstrained control policy, we trained a reinforcement learning agent to steer the network from three resistant attractor states to the target sensitive phenotype as in Table I.
-The resulting policy was exceptionally robust, achieving a 100% success rate over 10,000 evaluation trials.
-An analysis of the agent’s actions reveals a highly specific and hierarchical control strategy (Table II).
-The agent’s policy converged on targeting the MAP2K3/LOXL2 axis; its most frequent intervention was to flip MAP2K3, followed by flipping LOXL2.
-This demonstrates that while static network analysis pointed to JUN as a master regulator, the dynamic, goal-oriented learning process discovered a more efficient control pathway centered on the upstream kinase MAP2K3.
-TABLE II: Action frequencies of the trained reinforcement learning agent over 10,000 evaluation episodes.
-The data reveals a highly specific control policy that converges on flipping MAP2K3/LOXL2 as the primary intervention, identifying it as the optimal dynamic control target.
-Action Times Chosen
-Flip Gene ’MAP2K3’ 9037
-Flip Gene ’LOXL2’ 3366
-Flip Gene ’TAGLN’ 659
-Flip Gene ’RELA’ 483
-Flip Gene ’WNT5A’ 293
-Flip Gene ’JUN’ 183
-III-D Temporal Inhibition Reveals a “Hit-and-Run” Therapeutic Strategy
-To identify a clinically relevant, transient intervention strategy, we performed a comprehensive analysis simulating a temporary “priming” inhibition of key targets (JUN, LOXL2, and MAPK3) for varying durations.
-After each priming phase, a pre-trained RL agent was tasked with guiding the network to the therapy-sensitive state.
-The results, visualized in Fig.
-2, identified a highly effective and specific optimal strategy: a 4-step temporary inhibition of LOXL2, which achieved the highest success rate of 93.45%.
-The non-monotonic relationship between inhibition duration and success rate for both LOXL2 and JUN provides strong evidence for a “hit-and-run” mechanism, where a precisely timed intervention is more effective than a sustained one.
-A striking and consistent finding across all experiments was that the agent’s overwhelming post-priming policy was to “Do Nothing”.
-This validates our hypothesis that a brief, targeted intervention is sufficient to destabilize the resistant network, allowing the system’s own dynamics to complete the transition to a sensitive state.
-Figure 2: Efficacy of “hit-and-run” priming strategies.
-The plot shows the control success rate as a function of the inhibition duration for three key targets.
-A 4-step inhibition of LOXL2 emerges as the optimal transient strategy, significantly outperforming other interventions.
+III-C Temporal Inhibition Reveals a “Hit-and-Run” Therapeutic Strategy
+To identify a clinically relevant transient intervention, we simulated temporary “priming” inhibition of key targets (JUN, LOXL2, MAPK3) for varying durations, followed by RL-guided control to the therapy-sensitive state.
+The results (Fig.
+2) identified an optimal strategy: 4-step LOXL2 inhibition achieving 93.45% success rate.
+The non-monotonic relationship between inhibition duration and success for both LOXL2 and JUN demonstrates a “hit-and-run” mechanism where precisely timed intervention outperforms sustained targeting.
+Critically, the agent’s overwhelming post-priming policy was “Do Nothing,” validating that brief intervention destabilizes resistance, allowing natural dynamics to complete transition.
+Figure 2: Efficacy of “hit-and-run” priming strategies. A 4-step inhibition of LOXL2 emerges as the optimal transient strategy, significantly outperforming other interventions.
+III-D Mechanistic Policy Explanation via SHAP
+We employed SHAP analysis[24] to understand the agent’s control reasoning, providing mechanistic explanations for its strategy.
 Figure 3: Comparative SHAP analysis for the agent’s decision to flip MAP2K3 versus JUN from a resistant state.
-The plot shows the SHAP values for each gene’s contribution to the respective action’s logit, revealing a context-dependent policy.
-III-E Mechanistic Policy Explanation via SHAP
-To deconstruct the sophisticated policy learned by the reinforcement learning agent, we employed SHAP[22].
-This allowed us to move beyond observing the agent’s actions to understanding its underlying reasoning, providing mechanistic explanations for its control strategy.
-III-E1 Context-Dependent Intervention Choice
-We sought to understand why the agent prioritizes flipping MAP2K3 over other influential genes like JUN.
-A comparative SHAP analysis reveals a highly context-dependent policy (Fig.
+III-D1 Context-Dependent Intervention Choice
+SHAP analysis reveals why the agent prioritizes MAP2K3 over JUN through context-dependent logic (Fig.
 3).
-While high LOXL2 states push the agent towards flipping either MAP2K3 or JUN, the states of NRAS and TAGLN act as decisive signals: high NRAS encourages MAP2K3 targeting, while high TAGLN favors JUN intervention.
-This context-dependent strategy aligns with melanoma signaling biology.
-NRAS, mutated in 15-20% of melanomas [24], activates the MAPK cascade through MAP2K3 [25].
-The agent’s preference for MAP2K3 in NRAS-high contexts reflects understanding of this upstream-downstream relationship, intercepting oncogenic signals at a strategic chokepoint.
-Conversely, TAGLN drives cytoskeletal remodeling and epithelial-mesenchymal transition [26], prompting the agent to target JUN—a key AP-1 transcription factor [27]—to disrupt transcriptional programs driving resistance.
-This demonstrates that the agent discerns specific molecular contexts rather than identifying a generic resistant state, selecting the most effective pathway-specific intervention and mirroring precision oncology strategies [28].
-III-E2 Validation of the “Hit-and-Run” Mechanism
-Next, we used SHAP to explain why the 4-step inhibition of LOXL2 was more successful than the sub-optimal 2-step inhibition of JUN.
-We analyzed the network state after each priming strategy by measuring the average SHAP importance of each gene.
-The results show that the optimal LOXL2 priming is more effective because it does a better job of erasing the “vulnerability signature” that triggers the agent’s desire to intervene (Fig.
+While high LOXL2 states favor flipping either MAP2K3 or JUN, NRAS and TAGLN states determine the choice: high NRAS directs MAP2K3 targeting, while high TAGLN favors JUN intervention.
+This strategy aligns with melanoma biology.
+In NRAS-mutant contexts (15-20% of melanomas [26]), the agent targets MAP2K3 to intercept MAPK cascade activation [27] at a strategic chokepoint.
+Conversely, when TAGLN drives epithelial-mesenchymal transition [28], the agent targets JUN [29] to disrupt resistance-associated transcriptional programs.
+This demonstrates pathway-specific intervention selection mirroring precision oncology strategies [30].
+III-D2 Validation of the “Hit-and-Run” Mechanism
+SHAP analysis explains why 4-step LOXL2 inhibition outperformed 2-step JUN inhibition by measuring residual “vulnerability signatures” (Fig.
 4).
-After the LOXL2 inhibition, the residual SHAP values for key driver genes like NRAS and RELA are significantly lower than after the JUN inhibition, indicating the network has been pushed to a more stable, less-resistant state.
-Figure 4: Comparison of residual SHAP feature importance after the optimal (4-step LOXL2) versus a sub-optimal (2-step JUN) priming strategy.
-The more effective LOXL2 priming leaves a state with lower overall SHAP importance, requiring less subsequent intervention.
+LOXL2 priming more effectively erases intervention triggers, leaving significantly lower SHAP values for key drivers like NRAS and RELA, indicating the network reached a more stable, less-resistant state.
+Figure 4: Residual SHAP importance after optimal (4-step LOXL2) versus sub-optimal (2-step JUN) priming. Effective LOXL2 priming requires less subsequent intervention.
+LOXL2’s superior efficacy reflects its pleiotropic role beyond matrix crosslinking—it acts as transcriptional regulator promoting epithelial-mesenchymal transition and coordinates multiple resistance pathways including hypoxia response and immune evasion[31].
+This aligns with its position as a master coordinator of resistance-associated programs[32].
+III-D3 Dynamic Visualization of the Control Strategy
+Analysis of 10,000 simulated episodes reveals persistent “vulnerability signatures” driving the agent’s policy (Fig.
+5).
+The most frequent action was consistently “Do Nothing,” providing quantitative support for “hit-and-run” control where networks self-correct after initial priming.
+The heatmap shows powerful initial negative SHAP contributions from NRAS and RELA (-0.4 to -0.6), indicating strong intervention impetus.
+This signal persists at lower intensity (-0.2 to -0.4) throughout episodes, suggesting these genes exert background pressure the agent learns to ignore.
+Other genes contribute weaker, transient signals before becoming neutral.
+Figure 5: SHAP trajectory analysis showing persistent vulnerability signatures.
+“Do Nothing” was most frequent action, with NRAS/RELA negative signals that agent learns to ignore, validating the “hit-and-run” strategy.
 […]
-Figure 5: SHAP trajectory analysis explaining the agent’s policy.
-The most frequent action was consistently “Do Nothing,” represented by a dash (-) on the x-axis.
-Persistent negative SHAP values for NRAS and RELA reveal a vulnerability signature that the agent learns to ignore, validating the “hit-and-run” strategy.
-III-E3 Dynamic Visualization of the Control Strategy
-Finally, to visualize the agent’s reasoning, we analyzed the average of 10,000 simulated 15-step control episodes (Fig.
-5).
-The resulting SHAP trajectory heatmap reveals that the most frequent action at every step was “Do Nothing,” providing strong quantitative support for a “hit-and-run” control mechanism where the network self-corrects after an initial priming.
-The heatmap illustrates a persistent “vulnerability signature” that drives the agent’s policy.
-At the initial step, we observe a powerful negative SHAP contribution from the key driver genes NRAS and RELA, indicating a strong initial impetus for intervention.
-While this signal weakens after the first step, it does not disappear; instead, it persists at a consistent, lower intensity for the remainder of the episode.
-This suggests that even as the network evolves, these genes continue to exert a background pressure that the agent has learned to correctly ignore.
-Other genes, including LOXL2 and MAPK3, contribute weaker, transient negative signals in the early steps before becoming neutral.
-This temporal pattern of network self-organization following an initial perturbation aligns with emerging therapeutic paradigms in cancer treatment.
-The concept parallels successful clinical strategies such as induction chemotherapy in acute leukemias, where intensive initial treatment creates system instability that resolves into remission [34], and immune checkpoint blockade, where brief T-cell activation triggers sustained anti-tumor responses [35].
-The identification of NRAS and RELA as key, persistent vulnerability drivers reflects established understanding that RAS/MAPK hyperactivation and NF-$\kappa$B inflammatory signaling represent critical resistance mechanisms [36, 1], supporting the biological plausibility of the discovered intervention strategy.
 IV Conclusion
@@
 Our analysis revealed that resistance corresponds to a fundamental system-level transition from a plastic, multi-stable network to a rigid, canalized landscape driven by a JUN/LOXL2 regulatory axis that locks the system into a robustly resistant state.
-Most significantly, our reinforcement learning approach discovered a non-obvious ”hit-and-run” therapeutic strategy: a precisely timed, 4-step temporary inhibition of LOXL2 achieving over 93% success in silico.
+Most significantly, our reinforcement learning approach discovered a non-obvious “hit-and-run” therapeutic strategy: a precisely timed, 4-step temporary inhibition of LOXL2 achieving over 93% success in silico.
 Explainable AI analysis demonstrated that this transient intervention erases the molecular vulnerability signature stabilizing resistance, allowing the network’s intrinsic dynamics to complete the therapeutic transition.
-Our findings are based on an in silico model with inherent limitations, including its binary gene representation, focus on a core gene set, and the specific reward configuration for reinforcement learning.
-For instance, using a positive action cost instead of zero would penalize frequent interventions and could yield a different optimal policy.
-Consequently, the identified strategies should be interpreted as computational hypotheses that require rigorous experimental validation in cell lines and animal models before any clinical consideration.
-The discovery of time-dependent optimal strategies challenges traditional paradigms focused on sustained pathway inhibition.
-The integrated PBN-RL-XAI pipeline offers a generalizable approach for discovering novel control strategies in complex, treatment-resistant diseases, providing a powerful framework for future therapeutic development.
-V Code Availability
-All source code developed for the PBN inference, reinforcement learning, and subsequent analyses presented in this paper are publicly available in our GitHub repository at https://github.com/Liu-Zhonglin/pbn-melanoma-project.
+Our findings have important limitations.
+The binary gene representation oversimplifies continuous expression dynamics and misses intermediate states.
+Our 12-gene focus, while IPRES-motivated, excludes tissue-specific factors and broader regulatory networks.
+The PBN framework’s discrete time steps and Boolean logic may not capture biological stochasticity and temporal dynamics.
+The model lacks spatial considerations, metabolic constraints, and immune interactions critical for in vivo efficacy.
+Additionally, zero action costs in our RL formulation could yield different policies under alternative reward structures.
+This framework has only been tested on a single melanoma dataset, requiring validation across diverse cancer types and genetic backgrounds.
+Clinical translation requires systematic validation: experimental testing in cell lines and organoids to confirm LOXL2 timing effects, preclinical studies in xenograft models combining temporal inhibition with immunotherapy, and development of biomarkers for patient selection.
+Clinical trials must establish optimal timing protocols, design adaptive frameworks for schedule optimization, and carefully monitor safety and resistance emergence.
+Phase I studies should focus on pharmacodynamics and immune activation patterns.
+The discovery of time-dependent optimal strategies challenges traditional sustained inhibition paradigms.
+While requiring extensive validation, the integrated PBN-RL-XAI pipeline offers a generalizable computational approach for discovering novel control strategies in complex, treatment-resistant diseases, providing a powerful framework for future therapeutic development.
 Acknowledgment